Find the points on the graph of the function f(x)=x+1/x-1 where the slope of the tangent line is equal to -1/3 Determine the equations of the tangent and normal line.
1
Expert's answer
2020-07-08T17:53:25-0400

Equation of tangent line is:

"y-y_0=f'(x_0)(x-x_0)"

We have:

"f'(x_0)=-1\/3"

Then:

"f'(x)=\frac{x-1-x-1}{(x-1)^2}=-\frac{2}{(x-1)^2}"

"-\frac{2}{(x_0-1)^2}=-1/3"

"(x_0-1)^2=6"

"x_0=\sqrt{6}+1" or "x_0=-\sqrt{6}-1"

"y_0=\frac{\sqrt{6}+2}{\sqrt{6}}" or "y_0=\frac{\sqrt{6}}{\sqrt{6}+2}"

So, equations of tangent lines:

"y-\frac{\sqrt{6}+2}{\sqrt{6}}=-\frac{1}{3}(x-\sqrt{6}-1)"

or

"y-\frac{\sqrt{6}}{\sqrt{6}+2}=-\frac{1}{3}(x+\sqrt{6}+1)"


Equation of normal line is:

"y-y_0=-\frac{1}{f'(x_0)}(x-x_0)"

So:

"y-\frac{\sqrt{6}+2}{\sqrt{6}}=3(x-\sqrt{6}-1)"

or

"y-\frac{\sqrt{6}}{\sqrt{6}+2}=3(x+\sqrt{6}+1)"